# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v1

@@ Abstract.
We prove a product theorem for sublinear bilipschitz equivalences which generalizes the classical work of Kapovich, Kleiner and Leeb on quasiisometries between product spaces. Independently we develop a new tool, based on a theorem of Cornulier on sublinear bilipschitz equivalences between solvable Lie groups, to evaluate the distortion of certain subgroups in central extensions. This is useful to provide lower bounds on Dehn functions. Building on work of Cornulier and Tessera, we compute the Dehn functions of all simply connected solvable Lie groups of exponential growth up to dimension $5$. We employ our product theorem to distinguish up to quasiisometry certain families among these groups which share the same dimension, cone-dimension and Dehn function. Finally, using a theorem of Peng together with our computations of cone-dimensions and Dehn functions, we establish the quasiisometric rigidity of the rank-two, five-dimensional solvable Lie group acting simply transitively by isometries on the horosphere in general position inside the product of three real hyperbolic planes.
@@
The first author was partially funded by Israel Science Foundation grant ISF 2919/19. The second author was funded by the DFG 281869850 (RTG 2229). We gratefully acknowledge the financial support from the ConYS programme of the Karlsruhe House for Young Scientists at the KIT, and from the PEPS JCJC programme of the INSMI
@@ 1. Introduction
Our purpose here is to demonstrate that the reduction procedure, in the present case at the level of $\rho_{1}$, has applications to the quasiisometric classification in the class of Lie groups of exponential growth, which is disjoint from that of nilpotent groups. This stems from two facts, both relying on [Cor11]:
  • Since the group $\rho_{1}(G)$ in the class $(\mathcal{C}_{1})$ have less complicated structure than $G$, it can happen that it splits in a direct product while $G$ did not. Under certain assumptions on the factors, we can then apply a suitable generalization of the Kapovich-Kleiner-Leeb theorems on quasiisometries between products ([KKL98]), in order to rule out the existence of a $O(u)$-equivalence (hence of a quasiisometry) between $G$ and $G^{\prime}$ when $\rho_{1}(G)$ and $\rho_{1}(G^{\prime})$ split as non-isomorphic direct products. This is the strategy of Theorem A and Theorem C below.
  • The distortion of certain subgroups in central extensions $\widetilde{G}$ of $G$ can be estimated by comparing $\widetilde{G}$ and $\rho_{1}(\widetilde{G})$. In return, this gives estimates on the Dehn function of $G$, which allows one to progress in the quasiisometry classification. We develop our second tool, Proposition D, for that purpose.
@@ 1.A. Sublinear bilipschitz equivalence and products
[…]
@@
[…]
@@
Theorems A and C allow us to distinguish between several families of completely solvable groups up to quasiisometry.
We will summarize this contribution for groups of dimension $5$ in Corollary 5.22. We now give an example of this strategy in dimension $4$.
@@
The parameter $\alpha$ in the definition of $G_{3,5}^{\alpha}$ can actually be taken in $(-1,1)\setminus\{0\}$; the two limit cases $\alpha=-1$ and $\alpha=1$ give the groups $\mathbf{R}\times\mathrm{Sol}_{3}$ and $\mathbf{R}\times P$ respectively, where $P$ is as above. A direct inspection of the asymptotic cone suffice to tell $G_{4,9}^{0}$ apart from the groups $\mathbf{R}\times G_{3,5}^{\alpha}$ when $\alpha<0$ and from $\mathbf{R}\times\mathrm{Sol}_{3}$.
Keeping the goal of quasiisometric classification in mind, results like Theorem C should be complemented by other ones establishing that among the class of groups considered, the subset of those in the class $(\mathcal{C}_{1})$ should be quasiisometrically closed. Peng notably achieved this for some groups [Pen11b, Corollary 5.3.8] by measuring the divergence of vertical geodesics; in her setting the groups of class $(\mathcal{C}_{1})$ are characterized by pure exponential divergence (and Peng is able to prove that the quasiisometries respect vertical geodesics). However Peng’s technique and results usually cannot be applied to the groups we consider here, since they are not nondegenerate in Peng’s sense.
1.B. Distortion, Dehn function, and applications
The next results are about distortion of one-parameter subgroups and the Dehn functions of completely solvable Lie groups. The Dehn function is a notorious quasiisometry invariant of compactly presented (and especially connected Lie) groups.
Definition 1.9.
Let $G$ be a simply connected Lie group, and let $X$ be a nonzero element of its Lie algebra $\mathfrak{g}$. The distortion function of a one-parameter subgroup $L=\{\exp(tX)\}$ in $G$ is the growth type of the function
$\Delta_{L}^{G}(r)=\sup\{t:\exp tX\in B_{G}(r)\}.$
The proposition below interpolates between [Osi01] (which corresponds to the case when the group $S$ in the statement is nilpotent) and [Osi02]; we use both theorems of Osin in the proof. Although it is logically independent from the aforementioned results, it rests on a more elaborate version of Theorem 1.6, also due to Cornulier (See Theorem 3.7).
Proposition D (Evaluating distortion in completely solvable groups).
Let $S$ be a completely solvable group with Lie algebra $\mathfrak{s}$. Let $X\in\mathfrak{s}$ be nonzero, and let $c_{X}=\sup\{j\in\mathbf{Z}_{\geqslant 1}\cup\{\infty\}\colon X\in C^{j}\mathfrak{s}\}$, where $C^{1}\mathfrak{s}=\mathfrak{s}$ and $C^{j+1}\mathfrak{s}=[\mathfrak{s},C^{j}\mathfrak{s}]$ for all $j\geqslant 1$. Let $L$ be the one-parameter subgroup generated by $X$ in $S$.
  1. (1)
    If $c_{X}<\infty$, and if $X$ is in a Cartan subalgebra (e.g. if $X$ is a regular or central element) then there exists $\alpha>0$ such that
    $\Delta_{L}^{S}(r)\asymp r^{c_{X}}.$
  2. (2)
    [Osi02] If $c_{X}=\infty$ then $\Delta_{L}^{S}$ is exponential.
As an illustration of Proposition D, using the relation between the Dehn function and the distortion in a central extension (Proposition 4.2), we can distinguish up to quasiisometry the two completely solvable groups of dimension $4$ and whose asymptotic cones have dimension $3$.
Example 1.10.
The groups named $S=G_{4,3}$ in [Mub63] and $S^{\prime}=\mathbf{R}^{2}\times A_{2}$ are the only four dimensional, completely solvable Lie groups of cone dimension $3$. Their Lie algebras $\mathfrak{g}_{4,3}$ and $\mathbf{R}^{2}\times\mathfrak{a}_{2}$ are spanned by $e_{1},\ldots,e_{4}$, subject to the following nonzero brackets:
$\mathfrak{g}_{4,3}:\quad$ $[e_{4},e_{1}]=e_{1},[e_{4},e_{3}]=e_{2}.$
$\mathbf{R}^{2}\times\mathfrak{a}_{2}:\quad$ $[e_{4},e_{1}]=e_{1}.$
Let $(\omega_{1},\ldots,\omega_{4})$ denote the basis of $Z^{1}(\mathfrak{g}_{4,3},\mathbf{R})$ dual to $(e_{1},\ldots e_{4})$ in $\mathfrak{g}_{4,3}$. Then $[\omega_{2}\wedge\omega_{4}]\in H^{2}(\mathfrak{g}_{4,3},\mathbf{R})$ generates a central extension (isomorphic to the Lie algebra $\mathfrak{g}_{5,10}$ in Table 2) whose generator is cubically distorted in the corresponding Lie group by Proposition D. It follows that the Dehn function of $G_{4,3}$ is at least cubic. On the other hand, $S^{\prime}$ has a symmetric, hence a CAT(0) left invariant Riemannian metric, so that the Dehn function of $S^{\prime}$ is quadratic. So $S$ and $S^{\prime}$ are not quasiisometric.
Remark 1.11.
It turns out that in order to compute the Dehn functions of completely solvable groups of dimensions 4 and 5, it is enough to use the versions of Proposition  D and Proposition 4.2 for nilpotent groups, which are well known (See e.g. [BW97] and also [Pit97, Theorem 3.1] for a different formulation). In Appendix B we provide an example (of dimension 13) for which our version for solvable groups is actually necessary and improves on previously known lower bounds techniques. See discussion in Section 4.D.
Proposition E.
The simply connected solvable Lie groups of dimension less or equal five and exponential growth have Dehn function of growth type $n$, $n^{2}$, $n^{3}$, $n^{4}$ or $\exp(n)$ exactly, all given for the groups of dimension less than $5$ in Table 4 and for the groups of dimension $5$ in Table 5.
Remark 1.12.
The Dehn functions for all solvable Lie groups of dimension up to $5$ and polynomial growth follow from those of nilpotent groups of these dimensions, which were completely determined by Pittet [Pit97, Proposition 7.1]. We restrict our computations to the groups of exponential growth in this paper.
Finally, we use theorems of Peng [Pen11a, Pen11b] that we combine with our work on the quasiisometric classification to reach a rigidity result for the group $\operatorname{Sol}_{5}$ (also named $G_{5,33}^{-1,-1}$ in [Mub63]), which is the semidirect product $\mathbf{R}^{2}\ltimes\mathbf{R}^{3}$ with diagonal action of the $\mathbf{R}^{2}$ torus, such that the three weights $\varpi_{1},\varpi_{2},\varpi_{3}\in\operatorname{Hom}(\mathbf{R}^{2},\mathbf{R})$ are linearly independent and sum to zero.
Among all the left-invariant metrics on this group, one makes it isometric to a hypersurface
=$\mathcal{H}=\{(z_{1},z_{2},z_{3})\in\mathbb{H}^{2}\times\mathbb{H}^{2}\times\mathbb{H}^{2}\colon b_{1}(z_{1})+b_{2}(z_{2})+b_{3}(z_{3})=0\}$
where for $i=1,2,3$, $b_{i}$ is a horofunction only depending on the projection to the $i$th factor. As such, it is a higher-rank generalization of the three-dimensional group $\mathrm{Sol}_{3}$ (or SOL) which admits the same description as a horosphere in $\mathbb{H}^{2}\times\mathbb{H}^{2}$.
Proposition F (Propositions 6.56.8).
The following hold:
  1. (1)
    Let $G$ be a group of class $(\mathcal{C}_{0})$, quasiisometric to $\operatorname{Sol}_{5}$. Then $G$ is isomorphic to $\operatorname{Sol}_{5}$.
  2. (2)
    Let $\Gamma$ be a finitely generated group quasiisometric to $\operatorname{Sol}_{5}$. Then there is a finite-index subgroup $\Gamma_{0}$ in $\Gamma$ and a homomorphism $\Gamma_{0}\to\operatorname{Sol}_{5}$ with finite kernel and whose image is a lattice in $\operatorname{Sol}_{5}$.
The distinctive property of $\operatorname{Sol}_{5}$ used in the proof of Proposition F is that, unlike all the other completely solvable Lie groups of the same dimension and cone-dimension, it has an exactly quadratic Dehn function. The fact that its Dehn function is quadratic was proved by Druţu [Dru04, Theorem 1.1] and Leuzinger-Pittet [LP04, Corollary 2.1].
Remark 1.13.
Passing to the finite index subgroup $\Gamma_{0}$ of $\Gamma$ is necessary in the quasiisometric rigidity statement (2) of Proposition F above, as the following example shows. Let $K$ be a number field of degree $3$ with Galois group $\Sigma=\operatorname{Sym}_{3}$ and consider the group $\Lambda=\operatorname{PSL}(2,\mathcal{O}_{K})$, which embeds as a non-uniform $\mathbf{Q}$-rank one lattice in $X=\mathbb{H}^{2}\times\mathbb{H}^{2}\times\mathbb{H}^{2}$. The group $\Sigma$ operates by automorphisms on $\Lambda$; these extend as conjugations in $G=\operatorname{Isom}(X)$ by maps of $\mathbb{H}^{2}\times\mathbb{H}^{2}\times\mathbb{H}^{2}$ permuting the three factors. If $\mathcal{H}$ is a horosphere bounding a cusp in the quotient $\Gamma\backslash X$, then by [Pra73, Proposition 2.1(3)], $\Lambda\rtimes\Sigma$ intersects $\operatorname{Isom}(\mathcal{H})$ in a uniform lattice of the latter group; let us denote this lattice by $\Gamma$. One may write $\Gamma=\Gamma_{0}\rtimes\Sigma$, where $\Gamma_{0}$ is a lattice in $\operatorname{Sol}_{5}<\operatorname{Isom}(\mathcal{H})$. Thus $\Gamma$ is quasiisometric to $\operatorname{Sol}_{5}$. However, if there was $\phi\colon\Gamma\to\operatorname{Sol}_{5}$ a group homomorphism with finite kernel, then $\ker\phi$ would contain $\Sigma$ (which is torsion), but intersect $\Gamma_{0}$ trivially, since $\Gamma_{0}<\operatorname{Sol}_{5}$ has no non-trivial finite subgroup; it would follow that $\ker\phi=\Sigma$ and $\Gamma$ would split as a direct product, which is not the case.
1.C. Organization of the paper
@@
Section 3 introduces some of the theory of completely solvable Lie groups and the significance of SBE to this theory. Corollary B and Theorem C are proved in Section 3.C. We actually deduce both of these statements from Theorem 3.17, although Corollary B could be derived directly. Theorem 3.17 is stated in terms of diagonal Heintze groups, which are the Gromov-hyperbolic groups of class $(\mathcal{C}_{1})$ and thus are of coarse type I.
Section 4 is dedicated to the proof of the lower bounds estimates of Dehn functions. In Section 4.A we prove the distortion estimates stated in Proposition D, and in Section 4.B we apply them to obtain the lower bounds. Section 4.D compares our results with other lower bound methods. In Appendix B we give an example in which our tool improves on previously known techniques.
Sections 5 and 6 are devoted to mapping the contribution of our work towards the quasiisometric classification of low dimensional completely solvable Lie groups. We list these groups up to dimension $5$ and compute their image by $\rho_{1}$ (Tables 1 and 2) and their Dehn functions (Tables 4 and  5). Our computations are based on a list of criteria, mostly following [CT17], which are explained in Section 5.A. We explain our computations in Section 5.C. In Corollary 5.22 we show which new groups we are able to distinguish up to quasiisometry using Theorem A. Section 5.D elaborates on particular families of completely solvable Lie groups which exhibit interesting behavior. Finally in Section 6 we prove the quasiisometric rigidity of the group $G_{5,33}^{-1,-1}$ (Proposition F), building on the work of Peng [Pen11a][Pen11b] for the part concerning finitely generated groups.
1.D. Acknowledgements
We thank Yves Cornulier for a useful discussion and pointing us to Theorems 6.E.2 and 10.H.1 in [CT17].
@@ 2.A. Preliminaries
Let $X$ and $Y$ be metric spaces. After fixing $x_{0}\in X$ and $y_{0}\in Y$, we denote by ||$\cdot$|| ⋅ | the distance to the respective basepoints in $X$ and $Y$. We denote $|x_{1}|\vee|x_{2}|:=\max\{|x_{1}|,|x_{2}|\}$.
@@
We say that $u$ is admissible if it is nondecreasing and lim $\limsup u(2r)/u(r)<+\infty$. Often, we will also assume that $u$ is subadditive, that is, $u(r_{1}+r_{2})\leqslant u(r_{1})+u(r_{2})$ for all $r_{1},r_{2}\geqslant 0$. This implies admissibility.
@@ 2.B. Going through cones
If $\sigma$ goes to +$\infty$+ ∞ we call this an asymptotic cone. When it is not relevant to mention all the remaining data, we will denote an asymptotic cone of $X$ by $X_{\omega}$.
@@ 2.C. SBE Preserve Product Structure
[…]
@@
[…]
@@
Let $x,x^{\prime}$ be $i$-horizontal. Notice two trivial facts: a) $d_{X_{i}}\big{(}\pi_{i}(x),\pi_{i}(x^{\prime})\big{)}=d_{X}\big{(}x,x^{\prime})$, and b) the geodesic $[x,x^{\prime}]$ in $X$ joining $x$ and $x^{\prime}$ is contained in a leaf of $X_{i}$. In particular, for each $x^{\prime\prime}\in[x,x^{\prime}]$, $(x,x^{\prime\prime})$ and $(x^{\prime},x^{\prime\prime})$ are $i$-horizontal pairs.
@@
We push the above argument a bit further in order to show that any two $d_{0}$-separated $i$-horizontal pairs are $j$-compatible, regardless of which fiber of $X_{i}$ they are in. Let $(x,x^{\prime})$ and $(z,z^{\prime})$ be two $d_{0}$-separated $i$-horizontal pairs. Assume first that $x,x^{\prime},z,z^{\prime}\in X$ are pairwise $i$-horizontal, so they all lie in the same fiber. Since each $X_{i}$ can be assumed to have infinite diameter, we can find a point $p$ in the same $X_{i}$ fiber which is far enough so that $(p,w)$ is $d_{0}$-separated for all $w\in\{x,x^{\prime},z,z^{\prime}\}$. It is clear that the $j$-type of all points must be equal to that of $B\big{(}p,99d_{0}(p)\big{)}$.
This allows us to define the $j$-type of each $X_{i}$-fiber, defined by a choice of $\hat{x}\in\hat{X}:=\prod_{m\neq i}X_{m}$. To finish the proof we need to show that the $(i,j)$-type of $\hat{x}\in\hat{X}$ depends only on $(i,j)$ and not on the choice of $\hat{x}$. Indeed, fix $\hat{x},\hat{z}\in\prod_{m\neq i}X_{m}$, and $\gamma:\mathbf{R}_{\geq 0}\rightarrow X_{i}$ an infinite ray in $X_{i}$ (we assume $X_{i}$ is of infinite diameter and geodesic and so by an Arzela-Ascoli argument there exists an infinite geodesic ray).
@@
The $O(u)$-embeddings. We denote $\hat{X}:=\prod_{m=2}^{n}X_{m}$, and write points $x\in X$ as a pair $(z,\hat{x})\in X_{1}\times\hat{X}$. Fixing $\hat{x}\in\hat{X}$ we obtain a map $\phi_{\hat{x}}:X_{1}\rightarrow Y_{1}$ given by $z\mapsto\pi_{Y_{1}}\big{(}\phi(z,\hat{x})\big{)}$. We prove that these are all $(L_{1},u_{\hat{x}})$-bilipschitz embeddings, for $u_{\hat{x}}=u+d_{0}+(u+d_{0})(|\hat{x}|_{\hat{X}})\in O(u)$, where $L_{1}\geqslant 1$ and $d_{0}\in O(u)$ are given by Lemma 2.17. Indeed, let $z,z^{\prime}\in X_{1}$ be a two points such that the pair $(z,\hat{x}),(z^{\prime},\hat{x})$ is $d_{0}$-separated. By assumption on $X_{1}$ this pair is $1$-uncompressed so
@@
If the pair $(z,\hat{x}),(z^{\prime},\hat{x})$ is not $d_{0}$-separated, then subadditivity of $d_{0}$ gives $d(z,z^{\prime})=d_{X_{1}}(z,z^{\prime})\leqslant d_{0}(z\vee z^{\prime})+d_{0}(\hat{x})$. Recalling that $L_{1}^{-1}<1$, this gives a lower bound:
$L_{1}^{-1}d(z,z^{\prime})-d_{0}(z\vee z^{\prime})-d_{0}(\hat{x})\leqslant 0\leqslant d\Big{(}\pi_{Y_{1}}\big{(}\phi(z,\hat{x})\big{)},\pi_{Y_{1}}\big{(}\phi(z^{\prime},\hat{x})\big{)}\Big{)}.$
@@
$d\Big{(}\pi_{Y_{1}}\big{(}\phi(z,\hat{x})\big{)},\pi_{Y_{1}}\big{(}\phi(z^{\prime},\hat{x})\big{)}\Big{)}\leq d\big{(}\phi(z,\hat{x}),\phi(z^{\prime},\hat{x})\big{)}\leq Ld(z,z^{\prime})+u(z\vee z^{\prime})+u(\hat{x}).$
We conclude that $\phi_{\hat{x}}:X_{i}\rightarrow Y_{i}$ is an $(L,u_{\hat{x}})$-bilipschitz embedding.
@@
(Equalities are in $(Y_{\omega})_{1}$). On the other hand $d\big{(}\phi(x^{n}),(y^{n}_{1},\hat{0}_{Y})\big{)}\leqslant u(y^{n}_{1})$ and $\lim_{\omega}\frac{y^{n}_{1}}{\sigma_{n}}=0$, so $\big{(}\phi(x^{n})\big{)}_{n\in\mathbf{N}}=(y^{n}_{1},\hat{0}_{Y})_{n\in\mathbf{N}}$ (equality in $Y_{\omega}$). By definition $\phi_{\hat{0}}(x^{n}_{1})=\pi_{Y_{1}}\big{(}\phi(x^{n}_{1},\hat{0})\big{)}$. We conclude
@@
Sublinear Control. Finally, we show there is $v\in O(u)$ such that for $\hat{x},\hat{w}\in\hat{X}$ and $x_{1}\in X_{1}$, we have $d_{Y_{1}}\big{(}\phi_{\hat{x}}(x_{1}),\phi_{\hat{w}}(x_{1})\big{)}\leq v(|(\hat{x},x_{1})|,|(\hat{w},x_{1}|)\leq v(\hat{x}\vee\hat{w})+v(x_{1})$. For convenience, we show it in the case where $\hat{w}=\hat{0}$, which clearly implies the general case. Assume towards contradiction that there are sequences $x^{n}_{1}\in X_{1}$ and $\hat{x^{n}}\in\hat{X}$ with $\lim_{\omega}\frac{u(\hat{x^{n}},x^{n}_{1})}{d_{Y_{1}}\big{(}\phi_{\hat{x^{n}}}(x^{n}_{1}),\phi_{\hat{0}}(x^{n}_{1})\big{)}}=0$. Denote $x^{n}=(x^{n}_{1},\hat{x^{n}})$ and $\sigma_{n}:=d_{Y_{1}}\big{(}\phi_{\hat{x^{n}}}(x^{n}_{1}),\phi_{\hat{0}}(x^{n}_{1})\big{)}$. Consider the cones $X_{\omega}=\operatorname{Cone}(X,x^{n},\sigma_{n}),Y_{\omega}:=\operatorname{Cone}\big{(}Y,\phi(x^{n}),\sigma_{n}\big{)}$. From Lemma 2.7 we know that the map $F=\operatorname{Cone}(\phi)$ is an $L$-bilipschitz homeomorphism hence respects the product structure of $X_{\omega}$ and $Y_{\omega}$. We thus have $F_{1}:(X_{\omega})_{1}\rightarrow(Y_{\omega})_{1}$ an $L$-bilipschitz map such that for $x_{\omega}:=(x^{n})_{n\in\mathbf{N}},x_{\omega}^{0}=(x^{n}_{1},\hat{0})_{n\in\mathbf{N}}$ we have:
@@
[…]

### v2

@@ Abstract.
We prove a product theorem for sublinear bilipschitz equivalences which generalizes the classical work of Kapovich, Kleiner and Leeb on quasiisometries between product spaces. We employ our product theorem to distinguish up to quasiisometry certain families among these groups which share the same dimension, cone-dimension and Dehn function; actually we do this by distinguishing them up to sublinear bilipschitz equivalence, which is slightly stronger. Especially, we recover the fact, recently obtained by Bourdon and Rémy with different groups, that there exists uncountably many quasiisometry classes of indecomposable, non-unimodular, higher rank solvable Lie groups.
@@
The first author was partially funded by Israel Science Foundation grant ISF 2919/19 of Yair Glasner, and grant ISF 2990/21 of Ori Parzanchevski. The second author was funded by the DFG 281869850 (RTG 2229). We gratefully acknowledge the financial support from the ConYS programme of the Karlsruhe House for Young Scientists at the KIT, and from the PEPS JCJC programme of the INSMI
@@ 1. Introduction
Our purpose here is to demonstrate that the reduction procedure, in the present case at the level of $\rho_{1}$, has applications to the quasiisometric classification in the class of Lie groups of exponential growth, which is disjoint from that of nilpotent groups. This stems from the following fact, relying on [Cor11]. Since the group $\rho_{1}(G)$ in the class $(\mathcal{C}_{1})$ have less complicated structure than $G$, it can happen that it splits in a direct product while $G$ did not. Under certain assumptions on the factors, we can then apply a suitable generalization of the Kapovich-Kleiner-Leeb theorems on quasiisometries between products [KKL98], in order to rule out the existence of an $O(u)$-equivalence (hence of a quasiisometry) between $G$ and $G^{\prime}$ when $\rho_{1}(G)$ and $\rho_{1}(G^{\prime})$ split as non-isomorphic direct products. This is the strategy of Theorem A and Theorem C below.
@@ 1.A. Sublinear bilipschitz equivalence and products
$\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{i}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\phi_{i}}$
@@
$M\times\prod_{i=1}^{n}X_{i}$$N\times\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{\sigma(i)}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{\sigma(i)}}$$\scriptstyle{\phi_{i}}$
@@
Theorems A and C allow us to distinguish between several families of completely solvable groups up to quasiisometry. The following is an example of this strategy in dimension $4$.
@@
The parameter $\alpha$ in the definition of $G_{3,5}^{\alpha}$ can actually be taken in $(-1,1)\setminus\{0\}$; the two limit cases $\alpha=-1$ and $\alpha=1$ give the groups $\mathbf{R}\times\mathrm{Sol}_{3}$ and $\mathbf{R}\times P$ respectively, where $P$ is as above. A direct inspection of the asymptotic cone suffice to tell $G_{4,9}^{0}$ apart from the groups $\mathbf{R}\times G_{3,5}^{\alpha}$ when $\alpha<0$ and from $\mathbf{R}\times\mathrm{Sol}_{3}$. As the group $\mathbf{R}\times P$ bears a symmetric metric with a Euclidean factor, one can describe the finitely generated groups quasiisometric to $\mathbf{R}\times P$ [KL01]; however, due to the presence of the Euclidean factor we do not know whether similar methods can describe the Lie groups quasiisometric to $\mathbf{R}\times P$; we raise this question in Appendix A.
For completely solvable groups, primary quasiisometry invariants are the cone dimension (that is, the covering dimension of the asymptotic cone) and the Dehn function. In a subsequent paper [GP], we use the wide range of tools developed in [CT17] by Cornulier and Tessera in order to determine the Dehn function of all completely solvable groups of exponential growth and dimension up to 5. In Corollary 4.2 below we will summarize the contribution of our strategy to the quasiisometric classification of these groups. More precisely, Corollary 4.2 lists the $5$-dimensional groups which share the same cone dimension and Dehn function, but for which Theorem C implies that they are nonetheless quasiisometrically distinct. We recover a fact already obtained recently (in a more specific form that we explain below) by Bourdon and Rémy [BR23].
Corollary D.
There exists uncountably many quasiisometry classes of indecomposable, non-unimodular, completely solvable Lie groups with quadratic Dehn function.
The family of pairwise non-isomorphic groups we consider in order to deduce this corollary are named $G_{5,19}^{1,\beta}$ with $\beta\in(0,+\infty)$ in [Mub63] while Bourdon and Rémy’s family is $G_{5,33}^{\alpha,1-\alpha}$ with $\alpha\in[1/2,+\infty)$ (it is indecomposable for $\alpha\neq 1$). All these groups have dimension 5 and cone dimension $2$, but the groups considered by Bourdon and Rémy have left-invariant proper CAT(0) metrics (a condition known to imply a quadratic isoperimetric inequality, see e.g. [Wen05]), while the groups in the $G_{5,19}^{1,\beta}$ family do not have such metrics, as can be checked using Azencott and Wilson’s criterion [AW76]. Note that Peng obtained uncountably many quasiisometry classes among the completely solvable groups [Pen11a, Pen11b]. The groups in the $G_{5,19}^{1,\beta}$ family fall outside of the CAT(0) groups and those considered by Peng, which are unimodular.
In order to formulate our final question, let us call a group of type NPC a completely solvable group with a CAT(0) left-invariant proper Riemannian metric, as characterized in [AW76]. Examples of groups of type NPC are provided by the $AN$ subgroups of (higher rank) semisimple Lie groups $G=KAN$, but this family is considerably more vast; in dimension $5$ already, it includes the uncountable family studied by Bourdon and Rémy and mentionned above.
Question 1.9.
In Theorem A, can one allow the factors $Y_{j}$ to be taken among non-abelian, indecomposable groups of type NPC?
The question can equally be raised in the weaker form, for the original version of the Kapovich-Kleiner-Leeb product theorem instead of Theorem A. Among all the examples in dimension $5$ that we reviewed in Section 4, none of Cornulier’s $\rho_{1}$-reductions (which are $O(\log)$-bilipschitz equivalences) could rule out a positive answer to Question 1.9.
1.B. Organization of the paper
@@
Section 3 introduces some of the theory of completely solvable Lie groups and the significance of sublinear bilipschitz equivalence (SBE) to this theory. Corollary B and Theorem C are proved in Section 3.C. We actually deduce both of these statements from Theorem 3.16, although Corollary B could be derived directly. Theorem 3.16 is stated in terms of diagonal Heintze groups, which are the Gromov-hyperbolic groups of class $(\mathcal{C}_{1})$ and thus are of coarse type I. Finally, in Section 4 we formulate and prove Corollary 4.2 about the families of $5$-dimensional Lie groups that can be distinguished up to quasiisometry using our product theorem; corollary D appears as a byproduct of this study.
1.C. Acknowledgements
We thank Yves Cornulier for a useful discussion. The second-named author thanks Marc Bourdon and Bertrand Rémy for sharing their results, and Antoine Velut for useful discussions.
@@ 2.A. Preliminaries
Let $X$ and $Y$ be metric spaces. After fixing $x_{0}\in X$ and $y_{0}\in Y$, we denote by $|\cdot|$ the distance to the respective basepoints in $X$ and $Y$. We denote $|x_{1}|\vee|x_{2}|:=\max\{|x_{1}|,|x_{2}|\}$.
@@
We say that $u$ is admissible if it is nondecreasing and $\limsup u(2r)/u(r)<+\infty$. Often, we will also assume that $u$ is subadditive, that is, $u(r_{1}+r_{2})\leqslant u(r_{1})+u(r_{2})$ for all $r_{1},r_{2}\geqslant 0$. This implies admissibility.
@@ 2.B. Going through cones
If $\sigma$ goes to $\infty$ we call this an asymptotic cone. When it is not relevant to mention all the remaining data, we will denote an asymptotic cone of $X$ by $X_{\omega}$.
@@ 2.C. SBE Preserve Product Structure
$\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{i}$$\scriptstyle{F}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{F_{i}}$
@@
    $\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{i}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\phi_{i}}$
@@
Let $x,x^{\prime}$ be $i$-horizontal. Notice two trivial facts: a) $d_{X_{i}}\big{(}\pi_{i}(x),\pi_{i}(x^{\prime})\big{)}=d_{X}\big{(}x,x^{\prime})$, and b) the geodesic $x,x^{\prime}]$ in $X$ joining $x$ and $x^{\prime}$ is contained in a leaf of $X_{i}$. In particular, for each $x^{\prime\prime}\in[x,x^{\prime}]$, $(x,x^{\prime\prime})$ and $(x^{\prime},x^{\prime\prime})$ are $i$-horizontal pairs.
@@
We push the above argument a bit further in order to show that any two $d_{0}$-separated $i$-horizontal pairs are $j$-compatible, regardless of which leaf of $X_{i}$ they are in. Let $(x,x^{\prime})$ and $(z,z^{\prime})$ be two $d_{0}$-separated $i$-horizontal pairs. Assume first that $x,x^{\prime},z,z^{\prime}\in X$ are pairwise $i$-horizontal, so they all lie in the same leaf. Since each $X_{i}$ can be assumed to have infinite diameter, we can find a point $p$ in the same $X_{i}$ leaf which is far enough so that $(p,w)$ is $d_{0}$-separated for all $w\in\{x,x^{\prime},z,z^{\prime}\}$. It is clear that the $j$-type of all points must be equal to that of $B\big{(}p,99d_{0}(p)\big{)}$.
This allows us to define the $j$-type of each $X_{i}$-leaf, defined by a choice of $\hat{x}\in\hat{X}:=\prod_{m\neq i}X_{m}$. To finish the proof we need to show that the $(i,j)$-type of $\hat{x}\in\hat{X}$ depends only on $(i,j)$ and not on the choice of $\hat{x}$. Indeed, fix $\hat{x},\hat{z}\in\prod_{m\neq i}X_{m}$, and $\gamma:\mathbf{R}_{\geq 0}\rightarrow X_{i}$ an infinite ray in $X_{i}$ (we assume $X_{i}$ is of infinite diameter and geodesic and so by an Arzela-Ascoli argument there exists an infinite geodesic ray).
@@
[…]
@@
[…]
@@
[…]
@@
(Equalities are in $(Y_{\omega})_{1}$). On the other hand $d\big{(}\phi(x^{n}),(y^{n}_{1},\hat{0}_{Y})\big{)}\leqslant u(y^{n}_{1})$ and $\lim_{\omega}\frac{y^{n}_{1}}{\sigma_{n}}=0$, so $\big{(}\phi(x^{n})\big{)}_{n\in\mathbf{N}}=(y^{n}_{1},\hat{0}_{Y})_{n\in\mathbf{N}}$ (equality in $Y_{\omega}$). By definition $)\phi_{\hat{0}}(x^{n}_{1})=\pi_{Y_{1}}\big{(}\phi(x^{n}_{1},\hat{0}$$\big{)}$. We conclude
@@
[…]
@@
$X_{\omega}$$Y_{\omega}$$\bar{X}_{\omega}$$\bar{Y}_{\omega}$$\scriptstyle{F}$$\scriptstyle{\bar{\pi}_{\omega}}$$\scriptstyle{\bar{\pi}_{\omega}}$$\scriptstyle{\bar{F}}$